# Orak: A Foundational Benchmark for Training and Evaluating LLM Agents on Diverse Video Games

### v1

@@ I.5 Result for Minecraft
Study on RL-based Fine-tuning. Although RL-based fine-tuning has demonstrated strong performance in many domains, such as mathematics and programming, we did not explore it in this study. Given the interactive nature of the Orak environments, it would be natural to derive dynamic, context-aware rewards from in-game feedback and apply RL fine-tuning methods such as DPO [89] or GRPO [90]. Unlike domains such as mathematics or programming, where problems typically have static correct solutions derived through logical reasoning, gameplay requires strategic reasoning that adapts dynamically to the actions of other agents. In games, the optimal action is often contingent on the evolving behavior of the player or opponents, reflecting the complex, interactive nature of multi-agent environments. This distinction is particularly relevant to real-world domains such as business, economics, and negotiation, where strategic decision-making is frequently modeled using game-theoretic frameworks [91]. Therefore, leveraging Orak as an environment for RL-based fine-tuning may significantly enhance an LLM’s capacity to reason strategically and operate effectively in multi-agent settings. This line of research holds promise for improving LLM performance in a broad range of real-world applications where understanding multi-agent dynamics is essential.
Support for Diverse Modalities. Orak supports evaluation of LLMs and VLMs, but it does not extend to other modalities often essential in real-world gameplay. One key example is sound. For instance, in Minecraft, players rely on zombie audio cues to avoid danger, while in Street Fighter III and StarCraft II, alert sounds indicate some specific attacks by the enemy. Also, beyond games in Orak, First-Person Shooter (FPS) games usually use gunfire sounds for spatial awareness, and horror games often rely on audio to signal the proximity of threats. Consequently, incorporating other modalities such as audio remains an open challenge, and benchmarking the performance of emerging Speech-Language Models [92, 93] in gaming scenarios could be an important step toward broader multimodal agent development.
Appendix R Broader Societal Impacts
Simulation games such as Minecraft and Stardew Valley offer rich environments where players can explore, mine resources, and craft items, enabling life-like simulations of human behavior. These games offer a valuable testbed for analyzing long-horizon behavior of LLM agents by systematically comparing gameplay trajectories of humans and LLM agents, which enables a rigorous assessment of whether LLMs exhibit human-like decision-making patterns. Moreover, introducing multiple agents into these environments allows for the study of emergent social behaviors among LLM agents. We believe such settings are particularly well-suited for precisely measuring the social impact of complex agent behaviors, offering valuable insights into the dynamics of LLM-based agents.

### v2

@@ I.5 Result for Minecraft
Study on RL-based Fine-tuning. Although RL-based fine-tuning has demonstrated strong performance in many domains, such as mathematics and programming, we did not explore it in this study. Given the interactive nature of the Orak environments, it would be natural to derive dynamic, context-aware rewards from in-game feedback and apply RL fine-tuning methods such as DPO (Rafailov et al., 2023) or GRPO (Shao et al., 2024). Unlike domains such as mathematics or programming, where problems typically have static correct solutions derived through logical reasoning, gameplay requires strategic reasoning that adapts dynamically to the actions of other agents. In games, the optimal action is often contingent on the evolving behavior of the player or opponents, reflecting the complex, interactive nature of multi-agent environments. This distinction is particularly relevant to real-world domains such as business, economics, and negotiation, where strategic decision-making is frequently modeled using game-theoretic frameworks (Zhang et al., 2024). Therefore, leveraging Orak as an environment for RL-based fine-tuning may significantly enhance an LLM’s capacity to reason strategically and operate effectively in multi-agent settings. This line of research holds promise for improving LLM performance in a broad range of real-world applications where understanding multi-agent dynamics is essential.
Support for Diverse Modalities. Orak supports evaluation of LLMs and VLMs, but it does not extend to other modalities often essential in real-world gameplay. One key example is sound. For instance, in Minecraft, players rely on zombie audio cues to avoid danger, while in Street Fighter III and StarCraft II, alert sounds indicate some specific attacks by the enemy. Also, beyond games in Orak, First-Person Shooter (FPS) games usually use gunfire sounds for spatial awareness, and horror games often rely on audio to signal the proximity of threats. Consequently, incorporating other modalities such as audio remains an open challenge, and benchmarking the performance of emerging Speech-Language Models (Chu et al., 2023, Cui et al., 2024) in gaming scenarios could be an important step toward broader multimodal agent development.
Societal Impact. Simulation games, Minecraft and Stardew Valley, offer rich environments where players can explore, mine resources, and craft items, enabling life-like simulations of human behavior. These games offer a valuable testbed for analyzing long-horizon behavior of LLM agents by systematically comparing gameplay trajectories of humans and LLM agents, which enables a rigorous assessment of whether LLMs exhibit human-like decision-making patterns. Introducing multiple agents into these environments allows for the study of emergent social behaviors among LLM agents. We believe such settings are particularly well-suited for precisely measuring the social impact of complex agent behaviors, offering valuable insights into the dynamics of LLM-based agents.